OpenStudy (anonymous):

What is the value of 4C2?

sam (.sam.):

\[_nC_r=\frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

4C2 = 4!/(2!*(4-2)!) = 4*3*2*1/(2*2) = 6

sam (.sam.):

Then just substitute n and r

OpenStudy (anonymous):

Oh okayyy.. I see how ya'l did that now.

OpenStudy (hba):

Use a calculator :)

OpenStudy (anonymous):

I don't know how to work this out?

OpenStudy (anonymous):

ok

sam (.sam.):

@montana1999 did you get it?

OpenStudy (hba):

@montana1999 Well see it's easy Here, n=4 r=2 You can use the formula. If you want me to teach you how to use a calculator i can do that :)

OpenStudy (anonymous):

How did ya'll become moderators?

OpenStudy (anonymous):

I didn't.

OpenStudy (amistre64):

i was borned a moderator ;)

OpenStudy (anonymous):

*born

OpenStudy (anonymous):

haha funny:D

OpenStudy (hba):

@montana1999 Become an Ambassdor instead :D

OpenStudy (hba):

Ambassador*